OpenStudy (anonymous):

derivative of (sqrt(t+2))^2

OpenStudy (anonymous):

first step is 2(sqrt(t+2))

OpenStudy (anonymous):

derivative of \[(\sqrt{t+2})^2\]

OpenStudy (anonymous):

actually that can be simplified to t +2

OpenStudy (anonymous):

before anything cant it

OpenStudy (agreene):

\[\left( \sqrt{t+2} \right)^{2}\]\[=t+2\] =1